OpenStudy (anonymous):

y=(x−5)^2−4 Identify the vertex, equation for the axis of symmetry, domain, range and x- and y-intercepts of the parabola. y=−2(x+8)^2+2 Identify the vertex, equation for the axis of symmetry, domain, range and x- and y-intercepts of the parabola.

OpenStudy (aum):

Compare it to the vertex form of the parabola y = a(x-h)^2 + k where (h,k) is the vertex and you can quickly identify the vertex.

OpenStudy (anonymous):

how about we get married? :DDD

OpenStudy (aum):

y=(x−5)^2−4 is an vertical parabola. Therefore the axis of symmetry will be a vertical line with the equation x = h where h ix the x-coordinate of the vertex found earlier.

OpenStudy (anonymous):

vertex = -b/2a domain is = x range = y x intercept is when y is = 0 y intercept is when x is = 0

OpenStudy (anonymous):

i got the vertex, axis of symmetry, domain, and range but i cant figure out the intercepts

OpenStudy (anonymous):

or the vertex can be (h,k) y = x - h)^2 + k

OpenStudy (aum):

To find x-intercept, set y = 0 and solve for x. To find y-intercept, set x = 0 and solve for y.

OpenStudy (anonymous):

you have this equation y=(x−5)^2−4 if you want to find the y intercept you just have to plug in the zero for the x y=((0)-5)^2 - 4 if you want to find the x intercept you just have to plug in thr zero for the y (0)=(x-5)^2-4

OpenStudy (anonymous):

so y=21 but for im not exactly getting a real answer for x

OpenStudy (anonymous):

i got x^2-10x-21=0

OpenStudy (aum):

(0)=(x-5)^2-4 4 = (x-5)^2 take square root +/- 2 = x-5 +2 = x - 5 gives x = 7 -2 = x - 5 gives x = 3

OpenStudy (anonymous):

you can just do this 0=(x-5)^2 - 4 4 = (x-5)^2 sqrt(4) = x - 5 2 = x - 5 x =7

OpenStudy (aum):

y-intercept: 21 x-intercepts: 3, 7

OpenStudy (anonymous):

or what aum wrote

OpenStudy (anonymous):

thank youu